OpenStudy (anonymous):

use the substitution to solve the following equation 4x-y-3z=-3 x-2y-2z=-3 x+y+3z=-2 a. -1 2 -1 b. 1 -2 1 c. -1 -2 1 d.1 -2 -1

OpenStudy (anonymous):

1) 4x - y -3z = -3 for -1 2 -1 Now LHS of equation = 4x - y - 3z = 4(-1) - (2) - 3(-1) = -4 -2 +3 = -6 + 3 = -3 = RHS Hence verified that LHS = RHS

OpenStudy (anonymous):

2) x - 2y - 2z = -3 for 1 -2 1 LHS = x - 2y - 2z = (1) - 2(-2) - 2(1) = 1 + 4 - 2 = 5 - 2 = 3 here LHS not equal to RHS (pls check if u have given values correct) u can do the rest similarly..........I hope it is clear now??

jhonyy9 (jhonyy9):

a. solutions are correct